A partnership has a commercial cleaning service with 35 employees. They are purchasing new vacuum cleaners, floor waxing equipment, and related equipment for $45,000. They estimate the equipment repairs and maintenance of the equipment will be $800/year. Additional supplies will be $300/year. Calculate the annual cost of the equipment using an ROI of 15%.

$15,762
$16,862
$17,400
$18,876
$18,999
Answer #1

Assuming life of machine is 4 year

Annaul cost f machine = Cost/PVIFA(15%,4years)

=45000/2.854978

=15762

Annaul cost of equipment = 15762+800+300 = 16862$